Knowing 1 inch = 2.54 cm, determine how many meters there are in a 640-yard race
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39616)

You still need two more pieces of information.


Answer by ikleyn(52776)

.
640 yards = 640*3 feet = 640*3*12 inches = 640*3*12*2.54 centimeters = 58521.6 centimeters = 585.216 meters.    ANSWER